# Soap Film Drainage Under Tunable Gravity Using a Centrifugal Thin Film Balance

### v1

@@ Abstract
Surface bubbles are an abundant source of aerosols, with important implications for climate processes. In this context, we investigate the stability and thinning dynamics of soap films under effective gravity fields. Experiments are performed using a centrifugal thin-film balance capable of generating accelerations from 0.2 up to 100 times standard gravity, combined with thin-film interferometry to obtain time-resolved thickness maps. Across all experimental conditions, the drainage dynamics are shown to be governed by capillary suction and marginal regeneration—a mechanism in which thick regions of the film are continuously replaced by thin film elements (TFEs) formed at the meniscus. We consistently recover a thickness ratio of 0.8–0.9 between the TFEs and the adjacent film, in agreement with previous observations under standard gravity. The measured thinning rates also follow the predicted scaling laws. We identified that gravity has three distinct effects: (i) it induces a strong stretching of the initial film, extending well beyond the linear-elastic regime; (ii) it controls the meniscus size, and thereby the amplitude of the capillary suction and the drainage rate; and (iii) it reveals an inertia-to-viscous transition in the motion of TFEs within the film. These results are supported by theoretical modeling and highlight the robustness of marginal regeneration and capillary-driven drainage under extreme gravity conditions.
@@ I Introduction
Surface bubbles are an abundant source of aerosols, whether above a glass of champagne [seon_effervescence_2017] or at the ocean surface [deike_mass_2022]. In the marine environment, the bursting of these bubbles is a primary mechanism for the generation of sea spray aerosols [woodcock_giant_1953, macintyre_flow_1972, wu_evidence_1981, leifer_secondary_2000, bird_daughter_2010], which play a critical role in atmospheric processes by influencing cloud formation, radiative balance, and the air–sea exchange of climate-relevant gases—making them key players in Earth’s climate system. The stability of surface bubbles is influenced by impurities or surfactants, which enable the thin liquid film to resist perturbations through Marangoni stresses. Nevertheless, the film gradually drains over time due to capillary suction, driven by a pressure difference between the film and the meniscus in contact with the liquid reservoir [lhuissier2012bursting, frostad_dynamic_2016, bhamla_interfacial_2017, lin_influence_2018, liu_dynamic_2018, miguet2021marginal]. This drainage results in a net flux from the film to the meniscus, causing progressive thinning and ultimately leading to film rupture as the structure becomes increasingly fragile.
@@
The same drainage process occurs in flat soap films [gochev2016chronicles]. In this case, the film is planar and connects a rigid frame or a liquid reservoir, terminating at a meniscus with curvature radius $r_{\rm m}$. This curvature induces a pressure deficit of $\gamma/r_{\rm m}$ in the meniscus relative to the flat film, where $\gamma$ is the surface tension. Such behavior is well documented in horizontal films, for instance in thin-film balances or equivalent geometries [sheludko_thin_1967, sonin_role_1993, joye_asymmetric_1994, velev_investigation_1995, cascao_pereira_bike-wheel_2001, yaminsky_stability_2010, chatzigiannakis_thin_2021, andrieux_microfluidic_2021]. As in the case of surface bubbles, drainage in flat films is not uniform but is influenced by the instability known as marginal regeneration, a mechanism originally proposed by Mysels et al. [mysels1959soap] and more recently extended by Gros et al. [gros2021marginal]. During this process, the initial film of thickness $h$ feeds into the meniscus, while thinner patches—thin film elements (TFEs) of thickness $h_{\rm TFE}<h$—are intermittently extracted, resulting in a net mass loss from the film. Despite the significance of this process, only a few studies have quantified the thickness contrast between the bulk film and the extracted elements, even though this parameter is essential to predictive models of film stability and drainage [lhuissier2012bursting, gros2021marginal, vigna_flowing_2025]. In vertical films, measurements near the lower horizontal meniscus report a typical ratio $h_{\rm TFE}/h\approx 0.8$ [nierstrasz1998marginal, yi_quantitative_2025] where these patches of TFEs detach following capillary-driven buoyancy [adami2015surface, couder1989hydrodynamics, plateau1873].
@@
The nucleation, growth, and detachment of TFEs is critical in the case of surface bubbles, where they invade the entire interface and control the overall thinning dynamics [lhuissier2012bursting, bhamla_placing_2016, frostad_dynamic_2016, miguet2021marginal]. In contrast, in vertical films, TFEs mostly nucleate along the lateral borders, then rise vertically and merge with the film. This leaves the central part of the film relatively smooth, and leads to a characteristic thinning behavior, where the rate is inversely proportional to film width [mysels1959soap, hudales1990marginal, berg2005experimental, seiwert2017velocity, monier_self-similar_2024], and thickness profiles follow universal self-similar forms [monier_self-similar_2024, auregan_drainage_2025]. In horizontal films, TFEs are also observed, although they remain localized near the meniscus [joye_asymmetric_1994, velev_investigation_1995, khristov_thin_2000, yaminsky_stability_2010, andrieux_microfluidic_2021, gros2021marginal, tregouet2021instability]. According to existing models, the linear flux of liquid exchanged from the film to the meniscus scales as [lhuissier2012bursting, vigna_flowing_2025, cantat_drainage_2025]:
@@
with $\eta$ the liquid viscosity and $k$ a dimensionless factor. While gravity does not appear explicitly in this expression, it influences the meniscus radius $r_{\rm m}$ by setting the curvature difference in surface bubbles [lhuissier2012bursting] or determining meniscus size in planar films [vigna_flowing_2025].
@@
In this study, we systematically vary the effective gravity to probe the stability and thinning dynamics of soap films across low and high gravity conditions. We test the robustness of existing models, particularly regarding the apparent universality of the thickness ratio, which remains constant across all gravities and liquid viscosities. A scaling law is established for the exchange flux, confirming that marginal regeneration governs the thinning process. More broadly, this work offers insight into how other body forces—magnetic [elias_magnetic_2005, moulton_reverse_2010, back_ferrofilm_2012, lalli_stability_2023] or electric fields [tsekov_streaming_2010, bonhomme_soft_2013, sett_experimental_2016, hussein_sheik_electroosmotic_2017]—may similarly influence drainage dynamics.
To this end, we developed a centrifugal thin film balance, which allows a soap film to spin about its axis of symmetry. It thereby offers a surface-bubble-like geometry without the complexity of curved interfaces or asymmetric borders typical of vertical films. This setup mimics gravity via centrifugal acceleration—achieving variations over more than two orders of magnitude—while suppressing lateral boundaries.
@@ II.2 Centrifugal thin film balance
The apparatus is depicted in Fig. 1(a). Soap films are formed on a horizontal circular frame with an internal radius $R=30$ mm by dipping it into the surfactant solution and withdrawing it over approximately one second, at a velocity of about $V\approx 100~\mathrm{mm\,s^{-1}}$. The frame is housed within a closed circular chamber, 90 mm in diameter and 40 mm in height, whose lower face is transparent. The transparent lower plate serves three purposes: (i) it enables optical access for imaging; (ii) it seals the chamber to maintain a humid environment, thereby suppressing evaporation which reduces the soap film lifetime [champougny2018influence, miguet2021life, boulogne2022measurement]; and (iii) it entrains the surrounding air in solid-body rotation with the film, ensuring both the film and the air rotate at angular speed $\omega$. Removing the plate disrupts the film velocity and acceleration conditions.
At time $t=0$, the entire system is set into rotation around the vertical axis passing through the center of the film, with an angular speed $\omega$ ranging from 8 to 170 rad.s-1 (motor range). This rotation induces a centrifugal acceleration $r\omega^{2}$, where $r$ is the radial distance from the rotation axis, effectively mimicking a tunable gravity-like force. The resulting acceleration at the film edge can be adjusted between 2 and 840 m·s-2, corresponding to approximately 0.2–85 times standard gravity, defining the effective gravity as $\tilde{g}=R\omega^{2}/g$.
@@ II.3 Film imaging and thickness measurements
The film is illuminated by a pulsed green laser operating at 532 nm (Litron NANO-L-125-15 PIV, 125 mJ per pulse, pulse width 8 ns). The laser is triggered by an optical fork that sends a signal once per rotation, synchronizing the emission of a pulse with each turn. This stroboscopic setup ensures that images are captured at the same angular position on each rotation, effectively freezing any orthoradial motion of the soap film. Images are recorded using a color Nova S12 camera. Both the laser and the camera are positioned at an angle of $70^{\circ}$ relative to the rotation axis and aligned such that the first reflection from the transparent plate is excluded from the camera’s field of view. Only rays that interfere within the film are collected, resulting in high-contrast interference images, as illustrated in Fig. 1. Once in rotation, the film exhibits concentric black and green interference fringes. At the periphery, the fringes appear rough and irregular due to the presence of TFEs generated along the meniscus. As shown below, these TFEs progressively invade the film, increasingly disturbing the interference pattern. Each fringe corresponds to an isothickness contour, with the thickness difference between a green and a black fringe of 140 nm. To determine the absolute thickness profile, a reference is required; this is taken as the last visible green fringe before the formation of the common black film at the center of the film [gochev2016chronicles]. At a given time, the film is thinner at its center, with the thickness gradually increasing toward the periphery.
@@ III.1 Typical dynamics
Upon initiating rotation, the soap film exhibits two distinct drainage Regimes following a short transient phase lasting approximately one second.
@@ III.2 Initial thickness profiles in Regime 1
In Fig. 3, we present the thickness profiles of the central zone in Regime 1. As noted above, these profiles remain essentially unchanged over time, except that the extent of the central zone gradually decreases. We therefore focus on the initial profile, obtained as soon as reliable thickness measurements become available.
@@
We tested several models to understand the initial thickness profiles.
Assuming incompressible interfaces [cantat_drainage_2025], the internal flow in this geometry reduces to a Poiseuille flow, and the centrifugal force leads to a film of uniform thickness thinning as [emslie_flow_1958]:
$h(t)=\frac{h_{0}}{\sqrt{1+\frac{\rho\omega^{2}h_{0}^{2}}{12\eta}t}}\approx\frac{\sqrt{12\eta}}{\sqrt{\rho}\omega}t^{-1/2},$ (2)
once $h$ becomes smaller than the initial thickness prior to centrifugation, $h_{0}$. Using typical values $\rho=1000$ kg·m-3, $\eta=1$ mPa·s, and $\omega=100$ rad·s-1, the film thickness reaches about 1 µm after approximately $10^{3}$ s. This timescale is much longer than the duration over which the central region is observed, indicating that internal flow can be neglected—consistent with the fact that the profiles do not show a strong dependence on liquid viscosity.
As a consequence, film stretching due to interfacial elasticity appears to be the most plausible mechanism [cantat_drainage_2025]. The observed profiles are therefore compared with a theoretical model based on the following assumptions. Since these profiles are not yet populated with TFEs, we assume they result from the stretching of an initially uniform film formed prior to centrifugation. Following the approach of Pasquet et al. [pasquet_thickness_2023] for vertical films under standard gravity, we assume that the film has an initial uniform thickness $h_{\rm 0}$ and surface tension $\gamma_{\rm 0}$ before rotation begins. Once the film is set into rotation, it is stretched by centrifugal forces, which are balanced by surface tension gradients [couder1989hydrodynamics, de2001young], leading to
@@
Because the initial surfactant concentration is above the critical micelle concentration (cmc), the film elasticity is assumed to be primarily associated with the presence of surface-active impurities [cantat_drainage_2025]. As the film stretches and thins, the local interfacial concentration of these impurities decreases. The surface tension can be approximated by $\gamma=\gamma_{\rm 0}+E\frac{h_{\rm 0}-h}{h_{\rm 0}}$, where $E$ denotes the interface elasticity [poryles2022non, pasquet_thickness_2023]. Combining these relations yields the following differential equation for the thickness $h(r)$:
@@
Eq. (5) predicts a sharp increase in thickness due to the combined exponential and quadratic dependencies. This prediction does not capture the experimentally observed irregular increase of $h(r)$ with $r$. Even though the theoretical approach assumes a constant interfacial elasticity, we can use Eq. (4) to infer an apparent local elasticity, $\tilde{E}(r)=\frac{\rho h_{\rm 0}\omega^{2}}{2}\,\frac{r}{\frac{d}{dr}\left(\ln h(r)\right)}$. To evaluate this expression, we estimate the initial thickness before rotation using Frankel’s law [champougny2015surfactant], $h_{\rm 0}\simeq 2\kappa^{-1}\left(\frac{\eta V}{\gamma_{\rm 0}}\right)^{2/3}$, with $\kappa^{-1}\approx 2~\mathrm{mm}$ the capillary length and $V\approx 100~\mathrm{mm\,s^{-1}}$ the characteristic pulling speed used to form the film. This yields $h_{\rm 0}\approx 80~\mu\mathrm{m}$.
In the inset of Fig. 3(a), $\tilde{E}(r)$ is shown as a function of rotation speed. In all cases, $\tilde{E}$ varies significantly across the radius—by up to three orders of magnitude—indicating that it cannot be considered constant, or that the film thickness before rotation is not perfectly uniform. As shown in the inset of Fig. 3(b), similar behavior is observed when varying viscosity. However, $\tilde{E}$ remains roughly constant near the film center and at the outer boundary. At the center, $\tilde{E}$ ranges between $10^{-7}$ and $10^{-6}\,\mathrm{N\,m^{-1}}$, more than three orders of magnitude smaller than the values reported by Poryles et al. [poryles2022non] for the same surfactant concentration (2.4 times above the cmc). At the outer edge, $\tilde{E}$ reaches a plateau whose value increases with rotation speed from $10^{-5}$ to slightly above $10^{-4}\,\mathrm{N\,m^{-1}}$. These values are still smaller than those typically reported in the literature, and their dependence on $\omega$ is inconsistent with elasticity governed by surface-active impurities. This suggests that an alternative description of the film elasticity is required in this geometry.
Thinning at the onset of rotation—prior to the development of marginal regeneration—cannot be explained by internal flow or by film stretching within the framework of current models. This calls for an improved description of film elasticity to account for this stage. However, in the following, a detailed characterization of the central zone thickness in Regime 1 is not required to understand the mechanisms of marginal regeneration, which govern film thinning at later times.
@@
Figure 4: Thickness profiles for the experiment shown in Fig. 1. (a) Profiles at different times with min and max fringe intensities. Circles denote the central region, which maintains a steady thickness (mean $\pm$ standard deviation over time), while squares correspond to the peripheral region. Dashed lines are guides to the eye. (b) Temporal evolution of $h_{\rm R}(t)$, $h_{\rm 0}(t)$, and $h_{\rm b}(t)$ on a log–log scale. Data are shown for every other interference fringe for clarity. The time $t^{\star}$ marks the full erosion of the central zone, defined by $h_{\rm b}(t^{\star})=h_{\rm 0}(t^{\star})$. (c) $h_{\rm b}(t)$ as a function of $h_{\rm R}(t)$, with open and closed symbols corresponding to times before and after $t^{\star}$, respectively. The dashed line shows a proportionality of 0.87.

### v2

# Soap Film Drainage Using a Centrifugal Thin Film Balance

@@ Abstract
Surface bubbles are an abundant source of aerosols, with important implications for climate processes. In this context, we investigate the stability and thinning dynamics of soap films under effective gravity fields. Experiments are performed using a centrifugal thin-film balance capable of generating accelerations from 0.2 up to 100 times standard gravity, combined with thin-film interferometry to obtain time-resolved thickness maps. Across all experimental conditions, the drainage dynamics are shown to be governed by capillary suction and marginal regeneration—a mechanism in which thick regions of the film are continuously replaced by thin film elements (TFEs) formed at the meniscus. We consistently recover a thickness ratio of 0.8–0.9 between the TFEs and the adjacent film, in agreement with previous observations under standard gravity. The measured thinning rates also follow the predicted scaling laws. We identified that effective gravity has three distinct effects: (i) it induces a strong stretching of the initial film, extending well beyond the linear-elastic regime; (ii) it controls the meniscus size, and thereby the amplitude of the capillary suction and the drainage rate; and (iii) it reveals an inertia-to-viscous transition in the motion of TFEs within the film. These results are supported by theoretical modeling and highlight the robustness of marginal regeneration and capillary-driven drainage under extreme gravity conditions.
@@ I Introduction
Surface bubbles are an abundant source of aerosols, whether above a glass of champagne [47] or at the ocean surface [18]. In the marine environment, the bursting of these bubbles is a primary mechanism for the generation of sea spray aerosols [55, 36, 56, 30, 8], which play a critical role in atmospheric processes by influencing cloud formation, radiative balance, and the air–sea exchange of climate-relevant gases—making them key players in Earth’s climate system. The stability of surface bubbles is influenced by impurities or surfactants, which enable the thin liquid film to resist perturbations through Marangoni stresses. Nevertheless, the film gradually drains over time due to capillary suction, driven by a pressure difference between the film and the meniscus in contact with the liquid reservoir [32, 21, 7, 6, 33, 34, 37]. This drainage results in a net flux from the film to the meniscus, causing progressive thinning and ultimately leading to film rupture as the structure becomes increasingly fragile.
@@
The same drainage process occurs in flat soap films [22]. In this case, the film is planar and connects a rigid frame or a liquid reservoir, terminating at a meniscus with curvature radius $r_{\rm m}$. This curvature induces a pressure deficit of $\gamma/r_{\rm m}$ in the meniscus relative to the flat film, where $\gamma$ is the surface tension. Such behavior is well documented in horizontal films, for instance in thin-film balances or equivalent geometries [49, 50, 26, 53, 12, 57, 15, 2]. As in the case of surface bubbles, drainage in flat films is not uniform but is influenced by the instability known as marginal regeneration, a mechanism originally proposed by Mysels et al. [41] and more recently extended by Gros et al. [23]. During this process, the initial film of thickness $h$ feeds into the meniscus, while thinner patches—thin film elements (TFEs) of thickness $h_{\rm TFE}<h$—are intermittently extracted, resulting in a net mass loss from the film. Despite the significance of this process, only a few studies have quantified the thickness contrast between the bulk film and the extracted elements, even though this parameter is essential to predictive models of film stability and drainage [32, 23, 54]. In vertical films, measurements near the lower horizontal meniscus report a typical ratio $h_{\rm TFE}/h\approx 0.8$ [42, 58] where these patches of TFEs detach following capillary-driven buoyancy [1, 16, 44].
@@
The nucleation, growth, and detachment of TFEs is critical in the case of surface bubbles, where they invade the entire interface and control the overall thinning dynamics [32, 7, 21, 37]. In contrast, in vertical films, TFEs mostly nucleate along the lateral borders, then rise vertically and merge with the film. This leaves the central part of the film relatively smooth, and leads to a characteristic thinning behavior, where the rate is inversely proportional to film width [41, 24, 5, 46, 39], and thickness profiles follow universal self-similar forms [39, 3]. In horizontal films, TFEs are also observed, although they remain localized near the meniscus [26, 53, 28, 57, 2, 23, 51]. According to existing models, the linear flux of liquid exchanged from the film to the meniscus scales as [32, 54, 11]:
@@
with $\eta$ the liquid viscosity and $k$ a dimensionless factor. While gravity does not appear explicitly in this expression, it influences the meniscus radius $r_{\rm m}$ by setting the curvature difference in surface bubbles [32] or determining meniscus size in planar films [54].
@@
In this study, we systematically vary the effective gravity and solution viscosity to probe the stability and thinning dynamics of soap films across low and high gravity conditions. Our objective is twofold. First, we test whether the thickness ratio $h_{\rm TFE}/h$ remains constant or depends on the effective gravity. Second, we examine the validity of Eq. (1) under varying effective gravity, assessing whether marginal regeneration remains the dominant thinning mechanism and whether effective gravity acts only indirectly by setting the meniscus radius $r_{\rm m}$. More broadly, this work provides insight into how other body forces—such as magnetic [19, 40, 4, 29] or electric fields [52, 9, 48, 25]—may similarly influence drainage dynamics by controlling the meniscus size, and hence the overall liquid exchange between the film and the meniscus.
To this end, we developed a centrifugal thin film balance, which allows a soap film to spin about its axis of symmetry. It thereby offers a surface-bubble-like geometry without the complexity of curved interfaces or asymmetric borders typical of vertical films. As in surface bubbles, the effective gravity is not spatially homogeneous. In surface bubbles, the relevant body force corresponds to the projection of the gravitational acceleration onto the film surface: its magnitude vanishes along the axis of symmetry and reaches a maximum at the contact between the film and the meniscus. A similar situation occurs in our setup, where the centrifugal acceleration acts as a body force that increases with distance from the axis of rotation. Importantly, its magnitude at the film–meniscus contact can be tuned over more than two orders of magnitude.
@@ II.2 Centrifugal thin film balance
The apparatus is depicted in Fig. 1(a). Soap films are formed on a horizontal circular frame with an internal radius $R=30$ mm by dipping it into the surfactant solution and withdrawing it over approximately one second, at a velocity of about $V\approx 100~\mathrm{mm\,s^{-1}}$. The frame is housed within a closed circular chamber, 90 mm in diameter and 40 mm in height, whose lower face is transparent. The transparent lower plate serves three purposes: (i) it enables optical access for imaging; (ii) it seals the chamber to maintain a humid environment, thereby suppressing evaporation which reduces the soap film lifetime [13, 38, 10]; and (iii) it entrains the surrounding air in solid-body rotation with the film, ensuring both the film and the air rotate at angular speed $\omega$. Removing the plate disrupts the film velocity and acceleration conditions. There is about 30 s between film formation and the onset of rotation.
At time $t=0$, the entire system is set into rotation around the vertical axis passing through the center of the film, with an angular speed $\omega$ ranging from 8 to 170 rad.s-1 (motor range). A transient of approximately 1 s is required for the system to reach rigid-body rotation. This rotation induces a centrifugal acceleration $r\omega^{2}$, where $r$ is the radial distance from the rotation axis, effectively mimicking a body force whose magnitude increases from the film center to the edge. The resulting acceleration at the film edge can be adjusted between 2 and 840 m·s-2, corresponding to approximately 0.2–85 times standard gravity, defining the effective gravity as $\tilde{g}=R\omega^{2}/g$.
@@ II.3 Film imaging and thickness measurements
The film is illuminated by a pulsed green laser operating at 532 nm (Litron NANO-L-125-15 PIV, 125 mJ per pulse, pulse width 8 ns). The laser is triggered by an optical fork that sends a signal once per rotation, synchronizing the emission of a pulse with each turn. This stroboscopic setup ensures that images are captured at the same angular position on each rotation, effectively freezing any orthoradial motion of the soap film. Images are recorded using a color Nova S12 camera. Both the laser and the camera are positioned at an angle of $70^{\circ}$ relative to the rotation axis and aligned such that the first reflection from the transparent plate is excluded from the camera’s field of view. Only rays that interfere within the film are collected, resulting in high-contrast interference images, as illustrated in Fig. 1. Once in rotation, the film exhibits concentric black and green interference fringes. At the periphery, the fringes appear rough and irregular due to the presence of TFEs generated along the meniscus. As shown below, these TFEs progressively invade the film, increasingly disturbing the interference pattern. Each fringe corresponds to an isothickness contour, with the thickness difference between a green and a black fringe of 140 nm. To determine the absolute thickness profile, a reference is required; this is taken as the last visible green fringe before the formation of the common black film at the center of the film [22, 60]. At a given time, the film is thinner at its center, with the thickness gradually increasing toward the periphery.
@@ III.1 Typical dynamics
Upon initiating rotation, the soap film exhibits two distinct drainage Regimes following the transient stage lasting approximately one second.
@@ III.2 Initial thickness profiles in Regime 1
In Fig. 3, we present the thickness profiles of the central zone in Regime 1. As noted above, these profiles remain essentially unchanged over time, except that the extent of the central zone gradually decreases. We therefore focus on the initial profile, obtained as soon as reliable thickness measurements become available, corresponding to the end of the transient stage.
@@
We tested several models to understand the initial thickness profiles of the central zone.
Assuming incompressible interfaces [11], the internal flow in this geometry reduces to a Poiseuille flow, and the centrifugal force leads to a film of uniform thickness thinning as [20]:
$h(t)=\frac{h_{i}}{\sqrt{1+\frac{\rho\omega^{2}h_{i}^{2}}{12\eta}(t-t_{i})}},$ (2)
[…]
As a consequence, film stretching due to interfacial elasticity appears to be the most plausible mechanism to describe the thickness profiles in the central region [11]. The observed profiles are therefore compared with a theoretical model based on the following assumptions. Since these profiles are not yet populated with TFEs, we assume they result from the stretching of an initially uniform film formed prior to centrifugation. Following the approach of Pasquet et al. [43] for vertical films under standard gravity, we assume that the film has an initial uniform thickness $h_{\rm 0}$ and surface tension $\gamma_{\rm 0}$ before rotation begins. Once the film is set into rotation, it is stretched by centrifugal forces, which are balanced by surface tension gradients [16, 17], leading to
@@
Because the initial surfactant concentration is above the critical micelle concentration (cmc), the film elasticity is assumed to be primarily associated with the presence of surface-active impurities [11]. As the film stretches and thins, the local interfacial concentration of these impurities decreases. The surface tension can be approximated by $\gamma=\gamma_{\rm 0}+E\frac{\varepsilon}{1+\varepsilon}$, where $E$ denotes the interfacial elasticity and $\varepsilon=(h_{0}-h)/h$ the interface extension [45, 43]. Combining these relations yields the following differential equation for the thickness $h(r)$:
@@
Eq. (5) predicts a sharp increase in thickness due to the combined exponential and quadratic dependencies. This prediction does not capture the experimentally observed irregular increase of $h(r)$ with $r$ (Fig. 3). Even though the theoretical approach assumes a constant interfacial elasticity, we can use Eq. (4) to infer an apparent local elasticity, $\tilde{E}(r)=\frac{\rho h_{\rm 0}\omega^{2}}{2}\,\frac{r}{\frac{d}{dr}\left(\ln h(r)\right)}$.
In the inset of Fig. 3(a), $\tilde{E}(r)$ is shown as a function of rotation speed for the solution without glycerol. In all cases, $\tilde{E}$ varies significantly across the radius—by up to three orders of magnitude—and the larger the rotation speed, the larger the variation. At the inner radius, values start around $10^{-5}$$10^{-4}$ N·m-1 and increase to $3\times 10^{-3}$$3\times 10^{-2}$ N·m-1, eventually reaching a plateau at the highest rotation speeds. These values can be compared to $10^{-3}$ N·m-1, the typical value reported by Poryles et al. [45] for the same surfactant solution and concentration (2.4 times above the cmc), associated with the presence of dodecanol molecules acting as insoluble surface-active species at the interface. While this value lies near the middle of the range observed at a given rotation speed, our data cannot be fully explained by this description of interfacial elasticity. One possible explanation is the very large interface extension observed in our experiments, estimated as $(h_{0}-h(r=0))/h(r=0)$, with values around 80—well beyond unity—leading to strongly nonlinear behavior and combined effects of both dodecanol and SDS molecules on the surface tension, even above the cmc, due to exchanges between the bulk and the interfaces [45]. As discussed previously, interface stretching may also couple with internal flow at the highest rotation speeds. A similar trend is observed when varying the viscosity at constant rotation speed (inset of Fig. 3(b)). These observations call for an improved description of film elasticity to account for interface stretching and film thinning during the transient stage.
[…]
@@
Figure 4: Thickness measurements for the experiment shown in Fig. 1. (a) Profiles at different times (colors). Squares denote values measured in the peripheral region (dashed lines are guides to the eye), while crosses indicate the mean thickness profile of the central zone, with horizontal bars representing the standard deviation of the position. The symbols corresponding to $h_{\rm R}$, $h_{0}$, and $h_{\rm b}$ are defined in the panel. (b) Temporal evolution of $h_{\rm R}(t)$, $h_{\rm 0}(t)$, and $h_{\rm b}(t)$ on a log–log scale. Data are shown for every other interference fringe for clarity. The time $t^{\star}$ marks the full erosion of the central zone, defined by $h_{\rm b}(t^{\star})=h_{\rm 0}(t^{\star})$. Same symbols as in (a). (c) $h_{\rm b}(t)$ as a function of $h_{\rm R}(t)$, with open and closed symbols corresponding to times before and after $t^{\star}$, respectively. The dashed line shows a proportionality of 0.87.